# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes (Experiment, Analysis & Benchmark Paper)

### v3

@@ Abstract.
We present an in-depth analysis of data discovery in data lakes, focusing on table augmentation for given machine learning tasks. We analyze alternative methods used in the three main steps: retrieving joinable tables, merging information, and predicting with the resultant table. As data lakes, the paper uses YADL (Yet Another Data Lake) – a novel dataset we developed as a tool for benchmarking this data discovery task – and Open Data US, a well-referenced real data lake. Through systematic exploration on both lakes, our study outlines the importance of accurately retrieving join candidates and the efficiency of simple merging methods. We report new insights on the benefits of existing solutions and on their limitations, aiming at guiding future research in this space.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/rcap107/retrieve-merge-predict.
1. Introduction
In our data-rich era, data lakes –large, loosely structured corpora of raw data– have emerged as central resources in both academia and industry (Nargesian et al., 2019). Within data lakes, data discovery is crucial, particularly through the integration of multiple tables. This process has gained significant attention (Fan et al., 2023a). In this work, we focus on a specific usage of data lakes that requires capturing only a fraction of the tables: augmenting a base table for a machine learning (ML) task, such as in the following example scenario.
@@ 1. Introduction
Alice, a data scientist, wants to predict the box office revenue of a movie: she has access to a table that contains information about movies (e.g., year of release, director, language, budget, …). While she can use the available data to train a model, finding more information on the subject would be beneficial; for example, joining the table about movies with a table on the cast of those movies might help, as some actors tend to appear in movies with higher revenue. Alice has access to some large repository of data, or to some search engine that she can query (Castelo et al., 2021; Google, 2023) to find additional tables. Her objective is therefore to find the tables that are most relevant to her task, to merge them with the original table and finally to use the improved table to build a model that predicts the box office revenue.
@@
Figure 1. The pipeline. Given a base table, the three main steps (Retrieve, Merge, Predict) augment it with the information from the lake to improve the prediction performance.
As depicted in Figure 1, research on this problem is scattered across three main steps that involve four tasks: (1) retrieving the tables that are joinable with the original table based on shared attributes (Fernandez et al., 2019; Zhu et al., 2019, 2016; Castelo et al., 2021; Fernandez et al., 2018; Dong et al., 2023), (2) merging the information by executing the joins that most improve the performance of the subsequent ML model (Esmailoghli et al., 2021; Liu et al., 2022; Galhotra et al., 2023; Deng et al., 2017a), (3) aggregating results in cases of one-to-many or many-to-many joins (Chepurko et al., 2020; Kanter and Veeramachaneni, 2015), (4) predicting with the model on the resulting table (Prokhorenkova et al., 2018; Grinsztajn et al., 2022). For each task a variety of solutions have been proposed, yet each method is evaluated within disparate datasets and for the specific task it solves, rather than as the component of a full pipeline. In the absence of a unified benchmark to compare and evaluate these methods, it becomes hard for users like Alice to select the most suitable combination of methods for their specific use case.
Our study aims to promote a deeper understanding of discovery within data lakes for an analytic task. Our dataset, YADL (Yet Another Data Lake) makes such analysis reproducible and extensible. We leverage YADL to systematically explore the underlying mechanisms of end-to-end learning pipelines within data lakes.
YADL is constructed on the YAGO knowledge base (Mahdisoltani et al., 2014) to serve as a controlled environment for testing various methods by providing a defined vocabulary that includes entity names, thus alleviating issues stemming from uncertain matches and other ambiguities. The synthesis procedure accepts parameters to specify the shape (arity and cardinality) and degree of redundancy of the resulting tables: this allows to test how robust discovery and assembly systems are in scenarios with different degrees of complexity. YADL tables include both numerical and categorical features.
We investigate the impact of the main steps in the end-to-end process of learning from data lakes: generating potential augmentation candidates, integrating them with the base table, and subsequently evaluating the effectiveness of this assembled pipeline. We also provide a range of base tables to enable a comprehensive study.
This study has two main objectives: firstly, determine which tasks of the pipeline affect the prediction performance the most; secondly, develop an equitable playing field that facilitates a holistic and comparative assessment of alternative solutions.
Our pipeline enables the analysis of various methods in the four tasks. Our findings suggest that the retrieval of accurate join candidates is the most important task (improving performance by 4.9% in median relative to the worst retrieval method), outweighing the importance of the ML model used (4.1% difference in median). Results show that relying on Jaccard Containment (i.e., the fraction of entities of the table to augment found in a join candidate) as a criterion is effective for retrieving join candidates that improve downstream prediction performance. How join candidates are used to augment the base table affects the prediction performance, although the larger difference comes from joining multiple candidate tables at the same time rather than only selecting the best candidate (a difference of 4.2%). Also, simpler aggregation methods outpace their complex counterparts in terms of speed (being up to 4.5 times faster) and the larger cost of more intricate methods does not yield proportionate gains in prediction performance.
@@
To validate the generality of our findings, we contrast the insights obtained from the analysis carried out via YADL with those obtained from Open Data US, a well-referenced real data lake (Nargesian et al., 2018; Galhotra et al., 2023), over 8 input base tables. This comparative approach ensures that YADL is representative enough to be a benchmark.
@@
  3. (3)
    We conduct an experimental study on 8 base tables over various data lakes, retrieval techniques, join selection methods, aggregation solutions, and ML models (Section 5). Results provide insights on the steps of pipeline learning from data lakes.
@@
2. Problem setting and related work
Problem setting. Consider a user training a ML model to predict some quantity. The corresponding quantity appears in the training data as a target column $Y$ of a base table $T$. Assume our user has also access to a large collection of tables (a data lake) $D=\{T_{1},T_{2},...,T_{m}\}$, some of which may contain additional information that can enrich the base table $T$. The data lake may be publicly available (Castelo et al., 2021; Mattmann et al., 2018; Srinivas et al., 2023; Hulsebos et al., 2023a; Nargesian et al., 2018) or private (e.g., corporate data). Each of these tables $T_{k}$ is a bi-dimensional collection of data organized in columns $C^{i}_{k}\in T_{k}$ that may include both categorical (names, codes etc.) and numerical data (price, revenue, tax rates etc.). While cross-table metadata such as foreign keys is not available in this setting, joining $T$ with some of the tables in $D$ would be beneficial for the target prediction task.
Given $T$ and $D$, a table $T_{k}\in D$ is considered to be joinable, or a join candidate, if at least one of the columns $C^{i}_{k}\in T_{k}$ has a non-empty intersection with one of the columns in table $T$: if ${}\exists Q\in T,~{}\exists C^{i}_{k}\in T_{k}~{}|~{}Q\cap C^{i}_{k}\neq\emptyset$.
The user is interested in optimizing the performance of the ML model on a collection of columns (or “features”) $X$ according to the quality of the prediction for the target $Y$ (e.g., the movie revenue). Columns in $X$ may come either from $T$, or from joined tables in $D$.
The different steps, and corresponding literature. In such a scenario, the user is likely to go over the following main steps: Join Candidate Retrieval, Join Candidate Merging (which includes the Join Selection and Aggregation tasks), and Prediction. Depending on the specific scenario, some of the tasks may be executed in a different order or not at all. In the following, we will drop Join from the names when clear from the context.
While we focus on the joinability problem, a related concept is that of table unionability, i.e., identifying which tables may be “appended” to $T$ to increase the number of rows without increasing the number of features (columns) (Khatiwada et al., 2023a; Fan et al., 2023b; Nargesian et al., 2018).
@@ 2. Problem setting and related work
Given a base table $T$ and a data lake $D$, the Candidate Retrieval task consists in discovering join candidates (“Retrieval” in Fig. 1). This task looks for tables that can be considered as “candidate joins” for the given base table. Depending on the scenario, different methods have been developed: some systems (Castelo et al., 2021; Mattmann et al., 2018) work as “search engines” of sort by querying an open data internal database; other methods (Zhu et al., 2019; Fernandez et al., 2019; Zhu et al., 2016; Fernandez et al., 2018; Khatiwada et al., 2022, 2023b) are instead instantiated “locally” on data that the user has direct access to. While integrating data is not a central focus of the ML literature, augmenting features via joins is recognized as key to ML (Paleyes et al., 2022; Kumar et al., 2016). Still, the blind addition of features may lead to diminishing or negative returns, which led to the development of systems that seek to integrate only correlated features (Esmailoghli et al., 2021; Santos et al., 2021). Alternative systems seek to augment $T$ by synthesising new features (Kanter and Veeramachaneni, 2015; Cvetkov-Iliev et al., 2023; Zhao and Castro Fernandez, 2022). An orthogonal research direction is concerned with the problem of search and augmentation while maintaining privacy (Huang et al., 2023a, b).
@@
Though implementations may vary, retrieval methods share some similarities: 1) they involve an offline phase to either index the data lake (Zhu et al., 2016; Fernandez et al., 2019), or train a model (Fan et al., 2023b; Dong et al., 2023), and 2) they rely on similarity metrics across columns to select candidate joins. Often, this metric is Jaccard Containment (JC), defined as
(1) Jaccard Containment := $\frac{|Q\cap C_{k}^{i}|}{|Q|},$
@@
where $Q\in T$ is a query column, $C_{k}^{i}\in D$ is a candidate column in table $T_{k}$, $|Q\cap C_{k}^{i}|$ is the cardinality of the intersection between the two sets, and $|Q|$ is the cardinality of the query set itself. Intuitively, if this ratio is high, then a large fraction of $Q$ is found in $C_{k}^{i}$, suggesting that the two columns should be joined. Exhaustively measuring the JC is expensive. Thus approximate methods have been developed for fast estimation of the containment (Zhu et al., 2016; Fernandez et al., 2019, 2018). Beyond Jaccard Containment, other metrics have been used, such as top-$k$ set similarity (Zhu et al., 2019), combinations of JC and embeddings-based similarities (Fan et al., 2023b), or other embeddings-based metrics (Cong et al., 2022; Habibi et al., 2020), as approximate nearest neighbor search (Dong et al., 2023) on language models. Methods based on Jaccard Similarity are vulnerable to typos, which led to the development of methods that perform fuzzy or semantic matching such as (Fan et al., 2023b; Mundra et al., 2023; Deng et al., 2017b); in this work, we focus on scenarios that do not involve fuzzy matching.
@@
Retrieval methods are typically designed to be used on large data lakes (potentially, up to millions of tables (Zhu et al., 2019; Fan et al., 2023b)) and be as exhaustive as possible to maximize recall. However, three issues arise. First, these methods do not assess the relevance of join candidates to the downstream task. While a large containment value may indicate that a join can be done, it provides no guarantee that this join is actually useful (Kumar et al., 2016). Second, the number of candidate joins could become too large for practical use. Manually identifying the best candidates is time-consuming, and performing all joins might be too expensive in terms of time/memory constraints (Santos et al., 2021). A user-defined threshold on the containment can filter the most promising joins, but deciding the correct threshold is problematic. Alternatively, it is possible to select only the top-$k$ candidates by containment. Third, Jaccard Containment does not take into account the cardinality of a column: in an extreme case, if a column $Q\in T$ contains only a single value, it would have perfect overlap with any column $C_{k}^{j}\in D$ that contains the same value. While this behavior may be desirable for some specific retrieval tasks, it further expands the number of candidates with high containment. These limitations have led to the development of methods tackling the Join Selection task. As executing joins can be expensive, different methods have been proposed to reduce this cost, from avoiding the join altogether (Kumar et al., 2016; Shah et al., 2017), to joining over sketches or coresets of the data (Wang et al., 2022; Santos et al., 2021). A more straightforward approach consists in testing the candidate joins and select only those that improve the prediction performance (Chepurko et al., 2020; Gong et al., 2023; Galhotra et al., 2023; Dong and Oyamada, 2022).
@@
After the set of join candidates has been retrieved, they get merged with the base table to augment it with their columns (Merge in Fig. 1).
The Join Selection task addresses the problem of identifying a subset of candidate joins that maximizes the prediction performance over a downstream task. Candidate selection is performed by profiling candidate joins according to various parameters (Galhotra et al., 2023; Flores et al., 2021), removing joins that are not useful through the application of rules (Kumar et al., 2016; Shah et al., 2017), or by executing each candidate join to find those that bring benefit (Esmailoghli et al., 2021; Chepurko et al., 2020; Liu et al., 2022; Galhotra et al., 2023). In this work, we focus on the latter strategy.
It is therefore necessary to address the difficulties that come from joining base table and candidates. We consider ML models that involve a specific set of samples (rows) with their features. Augmentation operations in this scenario enrich said features while keeping the original set of samples constants, thus requiring a left join. A complication with left joins is that, for any un-joined row in the left table, null values will be added in the new columns. Alternative methods based on Full Disjunction (Galindo-Legaria, 1994; Khatiwada et al., 2022) add irrelevant samples that should then be dropped.
Note that, while we split the Candidate Join Retrieval and Selection steps for clarity, there is no hard separation between them. Indeed, some methods cover multiple steps (Gong et al., 2023; Huang et al., 2023c).
It is not always possible to limit joins to one-to-one relations; often, we join on one-to-many or many-to-many relations. For example, to join a table about movies with one that contains movie ratings on column “movie title”, every movie with more than one rating is involved in a one-to-many relation, thus the content of all the rows with such relations gets duplicated. Sample duplication is problematic when the downstream task involves a ML or statistical analysis method, as such models assume that each row correspond to one sample with its features. Thus, the Aggregation task bridges the result of the join with the downstream methods, combining the information contained from a potentially large number of rows into one. More precisely, given a tuple $t$ in the base table $T$ that needs to be joined with $n$ tuples from a binary table ($A$, $B$) over $A$, the goal is to augment $t$ with attribute $B$ by selecting one value that represents the information from the $n$ joining tuples. How to select a representative value for the new attribute reminds traditional data integration problems such as truth discovery (Dong et al., 2009) or data fusion (Bleiholder and Naumann, 2009). In this spirit, Deep Feature Synthesis (DFS) (Kanter and Veeramachaneni, 2015) takes a set of tables and a join plan to recursively aggregate replicated instances using functions such as average, median, and mode.
@@
Finally, the integrated table is used to train a model to predict a target variable through simple methods such as linear regression/classification (Pedregosa et al., 2011), or more complex ones such as CatBoost (Prokhorenkova et al., 2018). Though we present the prediction task as distinct from the prior ones, an optimized pipeline may implement learning at intermediate sections of the pipeline to ensure the selection of the most appropriate candidate tables(Galhotra et al., 2023; Huang et al., 2023c).
3. Building Yet Another Data Lake

### v6

@@ Abstract
Machine-learning from a disparate set of tables, a data lake, requires assembling features by merging and aggregating tables. Data discovery can extend autoML to data tables by automating these steps. We present an in-depth analysis of such automated table augmentation for machine learning tasks, analyzing different methods for the three main steps: retrieving joinable tables, merging information, and predicting with the resultant table. We use two data lakes: Open Data US, a well-referenced real data lake, and a novel semi-synthetic dataset, YADL (Yet Another Data Lake), which we developed as a tool for benchmarking this data discovery task. Systematic exploration on both lakes outlines 1) the importance of accurately retrieving join candidates, 2) the efficiency of simple merging methods, and 3) the resilience of tree-based learners to noisy conditions. Our experimental environment is easily reproducible and based on open data, to foster more research on feature engineering, autoML, and learning in data lakes.
1 Introduction: learning on data lakes needs feature engineering
New tools keep facilitating data science with machine learning, partly automating it, as with autoML (He et al., 2021; Karmaker et al., 2021; Erickson et al., 2020; Feurer et al., 2022; Hutter et al., 2019). However, they typically take a single table as an input, while data scientists often start from a data lake: a loose corpora of tables (Nargesian et al., 2019; Spotify, 2020; Phan, 2023). Learning then requires assembling multiple tables, which needs either in-depth knowledge of the data or data discovery (Fan et al., 2023a; Hulsebos et al., 2024). Data assembly is typically studied in scenarios where the schema is known (Robinson et al., 2024; Wang et al., 2024): how columns of various tables are related and which can be merged, as in a relational database. However, data lakes often come without this information. In this work, we consider a data-lake situation where the meaningful merges are unknown and where only a fraction of the tables in the data lake are relevant for the given machine learning (ML) task. Consider the following example scenario:
@@ 1 Introduction: learning on data lakes needs feature engineering
Alice, a data scientist, wants to predict the rating given by users to a movie: she has access to a table with information about movies (e.g., year of release, director, language, budget, …). She wants to complete the table, finding more information on the subject beneficial to her task; for example, joining the table about movies with a table on the cast of those movies might help, as some actors tend to appear in movies with better ratings. Alice has access to some large repository of data, or to some search engine that she can query to find additional tables (Castelo et al., 2021; Google, 2023). However, she does not know the schema describing the relation between tables in the data repository. Her objective is to find the tables that are most relevant to her task, to merge them with the original table and finally to use the improved table to build a model that predicts the movie rating. She has limited resource (compute, RAM…), and thus cannot evaluate all potential tables.
@@
Figure 1: The evaluation pipeline. Given a base table, the three main steps (Retrieve, Merge, Predict) augment it with the information from the lake to improve the prediction performance.
Research on this literature is spread across the database and machine learning communities. The corresponding pipeline (Figure 1) uses retrieve, merge, and predict steps that involve four tasks: 1) retrieving tables that are joinable with the original table (Fernandez et al., 2019; Zhu et al., 2019; 2016; Castelo et al., 2021; Fernandez et al., 2018; Dong et al., 2023), 2) selecting which joins should be executed to improve the performance of the subsequent ML model (Esmailoghli et al., 2021; Liu et al., 2022; Galhotra et al., 2023; Deng et al., 2017a), 3) aggregating results in cases of one-to-many or many-to-many joins (Chepurko et al., 2020; Kanter & Veeramachaneni, 2015), 4) predicting with supervised learning on the resulting table, a tabular learning problem (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022; Grinsztajn et al., 2022). Elaborate methods have been designed for each of these tasks separately. Yet, how they contribute to the overall machine learning pipeline is often unclear. Prediction performance is not evaluated consistently in the database literature. Moreover, evaluation on data lakes is particularly challenging. First, there is no openly-available reference data suited for this purpose: both a data lake and tables with corresponding analytic tasks. Second, a solid evaluation must run the pipeline many times across loosely structured, messy ensembles of tables. State-of-the-art publications relevant to this pipeline seldom come with implementations robust and scalable enough for cross-validation loops on a data lake. Indeed, data discovery and assembly code can involve many complex elements, such as probabilistic data structures or large language models, which need careful software engineering to be production ready. The difficulty of operating those complex pipelines does beg the question of when the prediction gain is worth the operational cost. State-of-the-art research often leaves aside the question of complexity, which involves not only computational cost, but also preparation and maintenance burden (Sculley et al., 2015; Paleyes et al., 2022).
We study the full data-discovery for machine learning pipelines, with an eye on exploring the complexity gradient in a reproducible way. For this, we contribute Yet Another Data Lake (YADL), a dataset that enables systematical exploration of the aspects of the pipeline important for good prediction. Built from the YAGO knowledge base (Mahdisoltani et al., 2014), YADL provides a controlled environment to test methods. To analyze the discovery and merge pipelines across different degrees of complexity, YADL’s synthesis procedure allows to vary the shape (number of rows and columns) and degree of redundancy of the resulting tables.
We investigate the impact of the main steps on the end goal of learning from data lakes: generating potential augmentation candidates, integrating them with the base table, and subsequently evaluating learning on this assembled data. Considering both retrieving and combining tables in a data lake and the downstream learning task, our work bridges database search and machine learning, a topic of much interest (Fey et al., ; Robinson et al., 2024; Cvetkov-Iliev et al., 2023; Wang et al., 2024). It has two main objectives: first, determine which pipeline steps are most important for prediction performance, and second, develop an equitable playing field that facilitates a holistic assessment of learning on data lakes.
We analyze various methods in the four tasks, with an exhaustive empirical evaluation that required approximately 21 years or 189k CPU and GPU hours (Table 7). Our findings show that 1) automated selection of joins is noisy, creating very messy tables for which tree-based methods are more robust than deep learning models; 2) better retrieval of joins reduces the amount of noise and the fraction of null values; 3) relying on Jaccard containment (i.e., the fraction of entities of the table to augment found in a join candidate) as a criterion for retrieving joins improves downstream prediction performance; 4) finally, simpler, metric-based join-discovery methods outpace their complex counterparts in terms of speed and the larger cost of more elaborate methods does not yield proportionate gains in prediction performance.
@@
To validate the generality of our findings, we contrast the insights obtained from the analysis carried out via YADL with those obtained from Open Data US, a well-referenced real data lake (Nargesian et al., 2018; Galhotra et al., 2023): this comparative approach ensures that YADL is representative enough to be a benchmark. We consider 6 input base tables from the literature which pertain to multiple domains.
@@
  3. 3.
    We conduct an experimental study on 6 base tables over various data lakes, retrieval techniques, join selection methods, aggregation solutions, and ML models (Section 5). Results provide insights on the steps of pipeline learning from data lakes.
@@
2 Problem setting and related work
Problem setting. Consider a user training a ML model to predict some quantity. The corresponding quantity appears in the training data as a target column $Y$ of a base table $T$. Assume our user has also access to a large collection of tables (a data lake) $D=\{T_{1},T_{2},...,T_{m}\}$, some of which may contain additional information that can enrich the base table $T$. The data lake may be publicly available (Castelo et al., 2021; Mattmann et al., 2018; Srinivas et al., 2023; Hulsebos et al., 2023; Nargesian et al., 2018) or private (e.g., corporate data). Departing from many studies on learning on relational databases, where schemas are provided (Motl & Schulte, 2015; Fey et al., 2023; Robinson et al., 2024), we consider unstructured data lakes, i.e., data lakes without a schema to specify any PK-FK (Private Key - Foreign Key) relationship between the tables.
Each of these tables $T_{k}$ is a bi-dimensional collection of data organized in columns $C^{i}_{k}\in T_{k}$ that may include categorical (names, codes etc.), numerical data (price, revenue, tax rates etc.) and text (product descriptions, user reviews etc.). While cross-table metadata such as foreign keys is not available in this setting, joining $T$ with some of the tables in $D$ would be beneficial for the target prediction task.
Given $T$ and $D$, a table $T_{k}\in D$ is considered to be joinable, or a join candidate for $T$ on a query column $Q\in T$, if at least one of the columns $C^{i}_{k}\in T_{k}$ has a non-empty intersection with one of the columns in table $T$: if ${}\exists Q\in T,~{}\exists C^{i}_{k}\in T_{k}~{}|~{}Q\cap C^{i}_{k}\neq\emptyset$.
The user wants to optimize the performance of the ML model on a collection of columns (or “features”) $X$ according to the quality of the prediction for the target $Y$ (e.g., the movie rating). Columns in $X$ may come either from $T$, or from joined tables in $D$.
Different steps. In such a scenario, the user is likely to rely on the following main steps: Join Candidate Retrieval, Join Candidate Merging (which includes the Join Selection and Aggregation tasks), and Prediction. Depending on the specific scenario, some of the tasks may be executed in a different order or not at all. In the following, we will drop Join from the names when clear from the context.
Another related problem, which we do not study, is that of finding new samples rather than features (called “table unionability” in databases): identifying which tables may be “appended” to $T$ to increase the number of rows (Khatiwada et al., 2023a; Fan et al., 2023b; Nargesian et al., 2018).
@@ 2 Problem setting and related work
Given a base table $T$ and a data lake $D$, the Candidate Retrieval task consists in discovering join candidates (Retrieve in Figure 1). This task looks for tables that can be considered as “candidate joins” for the given base table. Different methods have been developed for different scenario: some “dataset search engines” tackle internet-scale crawls of tables (Castelo et al., 2021; Mattmann et al., 2018); other methods are instantiated on data stored locally (Zhu et al., 2019; Fernandez et al., 2019; Zhu et al., 2016; Fernandez et al., 2018; Khatiwada et al., 2022; 2023b).
@@
While integrating data is not a central focus of the ML literature, augmenting features via joins is recognized as key to ML (Paleyes et al., 2022; Kumar et al., 2016). Still, the blind addition of features may lead to diminishing or negative returns, which led to the development of systems that seek to integrate only correlated features (Esmailoghli et al., 2021; Santos et al., 2021). Alternative systems seek to augment $T$ by synthesising new features (Kanter & Veeramachaneni, 2015; Cvetkov-Iliev et al., 2023; Zhao & Castro Fernandez, 2022). Recent contributions focus on building graph representations of relational databases to leverage GNNs for learning across tables (Fey et al., ; Robinson et al., 2024; Wang et al., 2024). An orthogonal research direction is concerned with the problem of search and augmentation while maintaining privacy (Huang et al., 2023a; b).
@@
Though implementations may vary, retrieval methods used to find join candidates share some similarities: 1) they involve an offline phase to either index the data lake (Zhu et al., 2016; Fernandez et al., 2019), or train a model (Fan et al., 2023b; Dong et al., 2023), and 2) they rely on similarity metrics across columns to select candidate joins. Often, this metric is Jaccard Containment (JC), defined as
Jaccard Containment := $\frac{|Q\cap C_{k}^{i}|}{|Q|},$ (1)
@@
where $Q\in T$ is a query column in base table $T$, $C_{k}^{i}\in D$ is a candidate column in table $T_{k}$, $|Q\cap C_{k}^{i}|$ is the cardinality of the intersection between the two sets, and $|Q|$ is the cardinality of the query set itself. Intuitively, if this ratio is high, then a large fraction of $Q$ is found in $C_{k}^{i}$, suggesting that the two columns should be joined. Approximate methods can scale JC-based retrieval to very large databases (Zhu et al., 2016; Fernandez et al., 2019; 2018). Beyond JC, other metrics have been used, such as top-$k$ set similarity (Zhu et al., 2019), combinations of JC and embeddings-based similarities (Fan et al., 2023b), or other embeddings-based metrics (Cong et al., 2022; Habibi et al., 2020), such as approximate nearest neighbor search (Dong et al., 2023) on model embeddings. Methods based on Jaccard similarity are vulnerable to typos, which led to the development of methods that perform fuzzy or semantic matching such as Fan et al. (2023b); Mundra et al. (2023); Deng et al. (2017b); in this work, we focus on scenarios that do not involve fuzzy matching.
@@
Retrieval methods are often designed for large data lakes (up to millions of tables Zhu et al., 2019; Fan et al., 2023b) and to maximize recall. However, three issues arise. First, these methods do not assess the relevance of join candidates to the downstream task. While a large containment value may indicate that a join can be done, it provides no guarantee that this join is actually useful (Kumar et al., 2016). Second, the number of candidate joins could become too large for practical use. Manually identifying the best candidates is time-consuming, and performing all joins might be too expensive in terms of time/memory constraints (Santos et al., 2021). A user-defined threshold on the containment can filter out the least promising joins, but deciding the correct threshold is problematic; alternatively, it is possible to select only the top-$k$ candidates by containment. Third, Jaccard containment does not take into account the cardinality of a column: in an extreme case, if a column $Q\in T$ contains only a single value, it would have perfect overlap with any column $C_{k}^{j}\in D$ that contains the same value. While this behavior may be desirable for some specific retrieval tasks, it further expands the number of candidates with high containment. The presence of duplicate tables worsens each of these issues by introducing potential false positives.
To make retrieval readily useful, the Join Selection task identifies a subset of candidate joins that maximizes the prediction performance over a downstream task (Select in Figure 1). These methods use various strategies to add value to retrieval: profiling candidate joins according to various metrics (Galhotra et al., 2023; Flores et al., 2021), rules to remove joins that are not useful (Kumar et al., 2016; Shah et al., 2017), joining over sketches or coresets of the data (Wang et al., 2022; Santos et al., 2021), or executing each candidate join to find those that bring benefit (Esmailoghli et al., 2021; Chepurko et al., 2020; Liu et al., 2022; Galhotra et al., 2023; Gong et al., 2023; Dong & Oyamada, 2022). In this work, we focus on the latter, very popular, strategy. We consider ML models that operate on rows as samples, columns forming features. Augmentation enriches features while keeping the original set of samples, thus requiring a left join. A resulting challenging is that un-joined rows in the left table will lead to missing (null) values in the new columns.
We split Candidate Join Retrieval and Selection to tease out the retrieval metric from the benefit for the task: the first measures similarities between columns, while the second depends on the information in the foreign table.
@@
After the set of join candidates has been retrieved, they must be merged with the base table to augment it with their columns. Depending on the join selection strategy, aggregation may be executed during the selection process, or after the set of candidates has been chosen. In any case, it is paired with the selection procedure by the need of training the ML model on the augmented tables. To represent this coupling, we combine the two tasks in a single Merge step in Figure 1.
It is not always possible to limit joins to one-to-one relationships; often, we join on one-to-many or many-to-many relations. For example, to join a table about movies with one that contains movie ratings on column “movie title”, every movie with more than one rating is in a one-to-many relation, thus the content of all the rows in such relations gets duplicated. Sample duplication is problematic when the downstream task involves a ML or statistical analysis method, as each row corresponds to one sample. The Aggregation task bridges the result of the join with the downstream methods, combining the information contained from a potentially large number of rows into one (Aggregate in Figure 1). More precisely, given a tuple (row) $t$ in the base table $T$ that needs to be joined with $n$ tuples from a table ($A$, $B$) over $A$, the goal is to augment $t$ with attribute $B$ by selecting one value that represents the information from the $n$ joining tuples. How to select a representative value for the new attribute reminds data integration problems such as truth discovery (Dong et al., 2009) or data fusion (Bleiholder & Naumann, 2009). In this spirit, Deep Feature Synthesis (DFS) (Kanter & Veeramachaneni, 2015) takes a set of tables and a join plan to recursively aggregate replicated instances using functions such as average, median, and mode.
@@
Finally, the integrated table is used to train a model to Predict a target variable through supervised methods. Tabular machine learning is the subject of very active research, with well-established tree-based methods (Grinsztajn et al., 2022; Prokhorenkova et al., 2018) and rapidly progressing tabular deep learning (Holzmüller et al., 2024; Ye et al., 2024; McElfresh et al., 2023; Hollmann et al., 2022; Hegselmann et al., 2023; Arik & Pfister, 2021; Gorishniy et al., 2024). It is not possible in our study to explore all recent methods, in particular given that some incur compute costs incompatible with our extensive evaluation and the combinatorics of pipelines. We consider three broad range of methods: 1) linear regression/classification (as implemented in scikit-learn Pedregosa et al., 2011), 2) tree-based methods such as CatBoost (Prokhorenkova et al., 2018), 3) or deep learning methods well suited for tabular data, such as ResNet (Gorishniy et al., 2021) or RealMLP (Holzmüller et al., 2024). While we present the prediction task as distinct from the prior ones, some pipelines implement learning at intermediate sections of the workflow to ensure the selection of the most appropriate candidate tables (Galhotra et al., 2023; Huang et al., 2023c).
3 Building Yet Another Data Lake